OpenStudy (anonymous):

Write the range of the following relation in list form. {(6.2, 0.3), (7.3, 0.3), (8.4, 0.3), (9.5, 0.3)} Range = {0.3} {6.2, 7.3, 8.4, 9.5} {1}

OpenStudy (anonymous):

Range =0.3 only (image of first value is 0.3only)

OpenStudy (anonymous):

{6, 7, 8, 9} {√2, √3, √4, √5} {6, 7, 8, 9, √2, √3, √4, √5}

OpenStudy (anonymous):

\[range is \left\{ \sqrt{2},\sqrt{3},\sqrt{4},\sqrt{5}\right\}\]

OpenStudy (anonymous):

A physics student is conducting an experiment of a falling object. He drops a baseball from the top of a building and records the distance traveled for each second of elapsed time. If the distances recorded are 16 ft., 64 ft., 144 ft., 256 ft., and 400 ft. respectively for each second of time, write this relation in ordered-pair form, using time for the domain. (16, 1), (64, 2), (144, 3), (256, 4), (400, 5)} {(1, 16), (2, 64), (3, 144), (4, 256), (5, 400)} {(1, 2), (3, 4), (4, 16), (64, 144), (256, 400)}